[ The length of a rectangle is 6 cm more than three times the width. If the perimeter of the rectangle is 28 cm, what are the dimensions? ]
Answer by TimothyLamb(4379)

x = 3y + 6
2x + 2y = 28
---
put the system of linear equations into standard form
---
x - 3y = 6
2x + 2y = 28
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
x= 12
y= 2
---
answer:
length = 12 cm
width = 2 cm
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php